# Generating Shrinkers by Mean Curvature Flow

### v2

@@ 9. Limits of Shrinkers
For all sufficiently large $g$, we have shown that, if $\Sigma_{g}\cap B(0,R)$ is connected, then it has genus $2g$, which is impossible since $R<R_{g}$. Thus we may assume, for all $g$, that $\Sigma_{g}\cap B(0,r)$ is not connected. Now it has at least $2$ components and at most $3$ (since $\Sigma_{g}\cap\partial B(0,R)$ has $3$ components.) The number of components is odd (by symmetry) and thus it is $3$. But now (exactly as in the proof of Claim 4) each of those components converges smoothly to the disk $\overline{D(0,R)}$, so $\Sigma_{g}\cap B(0,R)$ has genus 00, contradicting $r_{g}<R$. This completes the proof in Case 1.
Case 2: $\Sigma^{+}$ is nonempty. By Claim 1, its boundary is a circle $C(\rho)$. By Claim 3, $\Sigma^{+}$ meets $\{z=0\}$ orthogonally, so
$\Sigma:=\Sigma^{+}\cup\Sigma^{-}\cup C(\rho)$
is a smoothly embedded, rotationally invariant shrinker $\Sigma$. Thus $\Sigma$ is a sphere or a cylinder.
@@
If $\Sigma$ is a sphere, then (for large $g$) $\Sigma_{g}\cap\partial B(0,R)$ consists of $m=2c+1$ simple closed curves, $c$ of which lie in $\{z>0\}$. Also, $\Sigma_{g}\cap Z^{+}$ consists $c+1$ points, $c$ of which are close to the origin and one of which is close to $\Sigma^{+}\cap Z$. Thus
@@
If, on the other, $\Sigma$ is a cylinder, then (for large $g$) $\Sigma_{g}\cap\partial B(0,R)$ consists of $m+2$ simple closed curves, $m$ of which lie near the circle $C(R)$ and two of which are near the circles of intersection of the cylinder $\Sigma$ and the sphere $\partial B(0,R)$. Note that $c+1$ of those curves lie in $\{z>0\}$. Furthermore, $\Sigma_{g}\cap Z$ consist of $m$ points, $c$ of which lie in $Z^{+}$.
@@
We have shown that (42) holds whether $\Sigma$ is a sphere or a cylinder. Since $\Sigma_{g}$ has genus $\leq 2g$, we see that $c=0$. This completes the proof of Theorem 57 in Case 2. ∎
@@
Suppose $0<R<\infty$. Suppose that $C$ is a cone that is smooth except at the origin, and suppose that $M$ is a smooth shrinker with boundary in $B(0,R)$ and that $M$ is smoothly asymptotic (with multiplicity one) to $C$. Suppose for each $i$ that $M_{i}$ is a smooth shrinker with boundary in $B(0,R)$. Suppose that $M_{i}$ converges smoothly (and with multiplicity one) to $M$ on compact subsets of $\mathbf{R}^{3}\setminus\overline{B(0,R)}$. Then the convergence $M_{i}\to M$ is uniform at infinity in the following sense: if $\lambda_{i}\to 0$, then $\lambda_{i}\Sigma_{i}$ converges smoothly (with multiplicity $1$) on compact subsets of $\mathbf{R}^{3}\setminus\{0\}$ to $C\setminus\{0\}$.
@@ 10. Connectivity
and such that $\Sigma_{n}\setminus B(0,r_{n})$ contains two connected components $C_{n}$ and $C_{n}^{\prime}$. (It may have more components.) Let $\Sigma_{n}^{\prime}$ be a surface in $\mathbf{R}^{3}\setminus B(0,r_{n})$ of least shrinker area that separates $C_{n}$ and $C_{n}^{\prime}$.
Then $\Sigma_{n}^{\prime}$ is a surface whose boundary is in $\partial B(0,r_{n})$. It is smoothly embedded away from the boundary, and its shrinker area is at most
@@
Now the $\Sigma_{n}^{\prime}$ are stable, so (after passing to a subsequence, they converge smoothly in $\mathbf{R}^{3}\setminus\{0\}$ to a stable shrinker $\Sigma^{\prime}$ (possibly with multiplicity).
By [gulliver-lawson], $\Sigma^{\prime}$ is smoothly embedded everywhere. But then $\Sigma^{\prime}$ is unstable, a contradiction.
@@ 11. Genus of Shrinkers
Suppose that $T\leq\operatorname{T_{\textnormal{fat}}}$ and that $p\in M(T)$. Let $\mathscr{S}$ be the collection of all shrinkers at $(p,T)$, i.e., the the set of all subsequential limits of
@@
(In other words, at each critical point $q$, both principal curvatures with respect to the inward pointing normal are $>1/|q|$.)
@@
Suppose not. Then there exist $\Sigma_{i}\in\mathscr{S}$, and critical points $q_{i}\in\Sigma_{i}$ such that $R_{i}:=|q_{i}|\to\infty$ and such that one of the principal curvatures at $q_{i}$ (with respect to the inward pointing normal) is $\leq 1/|q_{i}|$.
@@
The mean curvature of $\Sigma_{i}$ at $q_{i}$ is $-q_{i}/2$. Thus the mean curvature of $R_{i}\Sigma_{i}$ at $R_{i}q_{i}$ is $-q_{i}/(2|q_{i}|)$.
Now translate the flow (*11) in spacetime by $(-R_{i}q_{i},R_{i}^{2})$ to get
@@
is mcf, i.e., such that $\Sigma^{\prime}$ is a translator with velocity $-q$. Thus the mean curvature at each point of $\Sigma^{\prime}$ is $-q^{\perp}$.
@@
be tangent flow at $\infty$ to the flow (43). Thus, there are $t_{n}\to-\infty$ such that
@@
converges to $\Sigma^{\prime\prime}$. By [bamler-kleiner], the convergence is smooth with multiplicity one. Note that the mean curvature vector $H_{n}$ at each point of $\Sigma^{\prime}_{n}$ is $(-t_{n}q)^{\perp}$. Thus
$|H_{n}|=|\nu\cdot H_{n}|=|t_{n}|\,|q^{\perp}|,$
@@
$|q^{\perp}|=|t_{n}|^{-1}|H_{n}|.$
Letting $n\to\infty$, we see that all points $q$ of $\Sigma^{\prime\prime}$, $|q^{\perp}|=0$. That is, $\Sigma^{\prime}$ is translation-invariant in the $q$ direction. Thus $\Sigma^{\prime}$ is the shrinking cylinder, and therefore $\Sigma^{\prime}$ is the bowl soliton [chh].
But that is a contradiction. ∎
@@
By the lemma and by standard Morse Theory, each component of $M\setminus B(0,r)$ is diffeomorphic to a closed disk or to $S^{1}\times[0,1)$. (Here $B(0,r)$ is the open ball.) ∎
@@
for all sufficiently large $i$. Let $R$ be as in the Lemma, and let $r>R$ be such that $\partial B(0,r)$ intersects $\Sigma$ transversely.

### v3

@@ 9. Limits of Shrinkers
For all sufficiently large $g$, we have shown that, if $\Sigma_{g}\cap B(0,R)$ is connected, then it has genus $2g$, which is impossible since $R<R_{g}$. Thus we may assume, for all $g$, that $\Sigma_{g}\cap B(0,r)$ is not connected. Now it has at least $2$ components and at most $3$ (since $\Sigma_{g}\cap\partial B(0,R)$ has $3$ components.) The number of components is odd (by symmetry) and thus it is $3$. But now (exactly as in the proof of Claim 4) each of those components converges smoothly to the disk $\overline{D(0,R)}$, so $\Sigma_{g}\cap B(0,R)$ has genus 0, contradicting $r_{g}<R$. This completes the proof in Case 1.
Case 2: $\Sigma^{+}$ is nonempty. By Claim 1, its boundary is a circle $C(\rho)$. By Claim 3, $\Sigma^{+}$ meets $\{z=0\}$ orthogonally, so the surface
$\widetilde{\Sigma}:=\Sigma^{+}\cup\Sigma^{-}\cup C(\rho)$
is a smoothly embedded, rotationally invariant shrinker. Thus $\widetilde{\Sigma}$ is a sphere or a cylinder.
@@
If $\widetilde{\Sigma}$ is a sphere, then (for large $g$) $\Sigma_{g}\cap\partial B(0,R)$ consists of $m=2c+1$ simple closed curves, $c$ of which lie in $\{z>0\}$. Also, $\Sigma_{g}\cap Z^{+}$ consists $c+1$ points, $c$ of which are close to the origin and one of which is close to $\Sigma^{+}\cap Z$. Thus
@@
If, on the other hand, $\widetilde{\Sigma}$ is a cylinder, then (for large $g$) $\Sigma_{g}\cap\partial B(0,R)$ consists of $m+2$ simple closed curves, $m$ of which lie near the circle $C(R)$ and two of which are near the circles of intersection of the cylinder $\widetilde{\Sigma}$ and the sphere $\partial B(0,R)$. Note that $c+1$ of those curves lie in $\{z>0\}$. Furthermore, $\Sigma_{g}\cap Z$ consist of $m$ points, $c$ of which lie in $Z^{+}$.
@@
We have shown that (53) holds whether $\widetilde{\Sigma}$ is a sphere or a cylinder. Since $\Sigma_{g}$ has genus $\leq 2g$, we see that $c=0$. This completes the proof of Theorem 61 in Case 2. ∎
@@
Suppose $0<R<\infty$. Suppose that $C$ is a cone that is smooth except at the origin, and suppose that $M$ is a smooth shrinker with boundary in $B(0,R)$ and that $M$ is smoothly asymptotic (with multiplicity one) to $C$. Suppose for each $i$ that $M_{i}$ is a smooth shrinker with boundary in $B(0,R)$. Suppose that $M_{i}$ converges smoothly (and with multiplicity one) to $M$ on compact subsets of $\mathbf{R}^{3}\setminus\overline{B(0,R)}$. Then the convergence $M_{i}\to M$ is uniform at infinity in the following sense: if $\lambda_{i}\to 0$, then $\lambda_{i}M_{i}$ converges smoothly (with multiplicity $1$) on compact subsets of $\mathbf{R}^{3}\setminus\{0\}$ to $C\setminus\{0\}$.
@@ 10. Connectivity
and such that $\Sigma_{n}\setminus B(0,r_{n})$ contains two connected components $C_{n}$ and $C_{n}^{\prime}$. (It may have more components.) Let $\Sigma_{n}^{\prime}$ be a surface in $\mathbf{R}^{3}\setminus B(0,r_{n})$ of least shrinker area that separates $C_{n}$ and $C_{n}^{\prime}$. Thus $\Sigma_{n}^{\prime}$ is a stable minimal surface with respect to the shrinker metric.
Now $\Sigma_{n}^{\prime}$ is a surface whose boundary is in $\partial B(0,r_{n})$. It is smoothly embedded away from the boundary, and its shrinker area is at most
@@
Since the $\Sigma_{n}^{\prime}$ are stable, they converge smoothly (after passing to a subsequence) in $\mathbf{R}^{3}\setminus\{0\}$ to a stable shrinker $\Sigma^{\prime}$ (possibly with multiplicity).
By [gulliver-lawson], $\Sigma^{\prime}$ is smoothly embedded everywhere. But then $\Sigma^{\prime}$ is unstable, a contradiction. (Every smooth shrinker without boundary is an unstable minimal surface with respect to the shrinker metric.)
@@ Appendix A Genus of Shrinkers
Suppose that $T\leq\operatorname{T_{\textnormal{fat}}}$ and that $p\in M(T)$. Let $\mathscr{S}$ be the collection of all shrinkers at $(p,T)$, i.e., the set of all subsequential limits of
@@
(In other words, at each critical point $q$, both principal curvatures with respect to the inward pointing normal are >1/|q|>1/|q|.)
@@
Suppose not. Then there exist $\Sigma_{i}\in\mathscr{S}$, and critical points $q_{i}\in\Sigma_{i}$ such that $R_{i}:=|q_{i}|\to\infty$ and such that one of the principal curvatures $\kappa_{i}$ at $q_{i}$ (with respect to the inward pointing normal) satisfies:
(54) $\kappa_{i}\leq 1/|q_{i}|.$
@@
The mean curvature of $\Sigma_{i}$ at $q_{i}$ is $q_{i}/2$. Thus the mean curvature of $R_{i}\Sigma_{i}$ at $R_{i}q_{i}$ is $q_{i}/(2|q_{i}|)$.
Now translate the flow (*A) in spacetime by $(-R_{i}q_{i},R_{i}^{2})$ to get
@@
is a mean curvature flow, i.e., such that $\Sigma^{\prime}$ is a translator with velocity q-q. Here q-q is the limit of $q_{i}/(2|q_{i}|)$. Thus the mean curvature at each point of $\Sigma^{\prime}$ is $q^{\perp}$.
@@
be a tangent flow at $\infty$ to the flow (55). Thus, there are $t_{n}\to$-$\infty$ such that
@@
converges to $\Sigma^{\prime\prime}$. By [bamler-kleiner], the convergence is smooth with multiplicity one. Note that the mean curvature vector $H_{n}$ at each point of $\Sigma^{\prime}_{n}$ is $(-|t_{n}|^{1/2}q)^{\perp}$. Thus
$|H_{n}|=|\nu\cdot H_{n}|=|t_{n}|^{1/2}\,|q^{\perp}|,$
@@
$|q^{\perp}|=|t_{n}|^{-1/2}|H_{n}|.$
Letting $n\to\infty$, we see that at all points of $\Sigma^{\prime\prime}$, $|q^{\perp}|=0$. That is, $\Sigma^{\prime\prime}$ is translation-invariant in the $q$ direction. Thus $\Sigma^{\prime\prime}$ is the shrinking cylinder, and therefore $\Sigma^{\prime}$ is the bowl soliton [chh].
But that is a contradiction, since one of the principal curvatures of $\Sigma^{\prime}$ at 0 with respect to the direction q-q is $\leq 0$, by (54). ∎
@@
By Lemma 66 and standard Morse Theory, each component of $M\setminus B(0,r)$ is diffeomorphic to a closed disk or to $S^{1}\times[0,1)$. (Here $B(0,r)$ is the open ball.) ∎
@@
for all sufficiently large $i$. Let $R$ be as in Lemma 66, and let $r>R$ be such that $\partial B(0,r)$ intersects $\Sigma$ transversely.